# Learning Content-enhanced Mask Transformer for Domain Generalized Urban-Scene Segmentation

### v1

@@ Abstract
In this paper, we propose a Content-enhanced Mask TransFormer (CMFormer) for domain-generalized USSS. The main idea is to enhance the focus of the fundamental component, the mask attention mechanism, in Transformer segmentation models on content information. To achieve this, we introduce a novel content-enhanced mask attention mechanism. It learns mask queries from both the image feature and its down-sampled counterpart, as lower-resolution image features usually contain more robust content information and are less sensitive to style variations. These features are fused into a Transformer decoder and integrated into a multi-resolution content-enhanced mask attention learning scheme.
@@
Extensive experiments conducted on various domain-generalized urban-scene segmentation datasets demonstrate that the proposed CMFormer significantly outperforms existing CNN-based methods for domain-generalized semantic segmentation, achieving improvements of up to 14.00% in terms of mIoU (mean intersection over union). The source code for CMFormer will be made available at this repository.
1 Introduction
@@
Urban-scene semantic segmentation (USSS) is a challenging problem because of the large scene variations due to changing landscape, weather, and lighting conditions [44, 31, 10, 4]. Unreliable USSS can pose a significant risk to road users. Hence, domain generalization is essential for robust USSS [35, 16, 7]. In contrast to common domain generalization, domain generalized USSS requires special attention because the domain gap is mainly caused by large style variations whereas changes in semantics largely remain consistent (see Fig. 1 for an example).
@@
Existing approaches can be divided into two groups. One group focuses on the style de-coupling. This is usually achieved by a normalization [35, 16, 38] or whitening [36, 7, 57, 38] transformation. However, the de-coupling methodology falls short as the content is not learnt in a robust way. The other group is based on adverse domain training [65, 23, 66]. However, these methods usually do not particularly focus on urban styles and therefore their performance is limited.
In this paper, we focus on two main objectives for USSS: (1) a concise but expressive representation of the content, and (2) a robust module for style change handling. For the first goal, we propose to use Mask Transformer. Transformer segmentation models [47, 5, 51, 60, 61, 6] provide more expressive representations than CNN-based methods because of their mask attention mechanisms allowing to incorporate long-range dependencies [27, 28, 11], context [19], and saliency [5, 6]. For the second goal, mask-level learning is used which is less sensitive than pixel-level learning in handling style changes [34, 12, 52].
Figure 1: (a) Example of domain generalized USSS (top: source domain, rest: target domain); (b) The key challenge lies in the domain gap is mainly from the extremely-varied styles but the content is similar; (c) The proposed CMFormer with content-enhanced mask attention focus on: 1) learn an expressive content representation, 2) robust to style variation. SAW: the present state-of-the-art [38].
@@
To this end, a novel Content-enhanced Mask TransFormer (CMFormer) is proposed. Particularly, a content-enhanced mask attention mechanism is introduced, which takes the original image feature together with its down-sampled counterpart as input. The down-sampled features contain more content and are less sensitive to style variations. Both features are fused to learn the context. Then, the content-enhanced mask attention is extended to incorporate multi-resolution features.
@@
Large-scale experiments are conducted with various domain generalized USSS settings, i.e., trained on one dataset from [42, 43, 8, 33, 59] as the source domain, and validated on the rest of the four datasets as the unseen target domains. The experiments show that the proposed CMFormer achieves upto 14.00% mIoU improvement compared to the state-of-the-art (e.g., SAW [38], WildNet [23]). The proposed CMFormer also outperforms existing mask-level Transformer segmentation methods and the generic segmentation methods pre-trained by the foundation model (i.e., SAM [21], SegGPT [54]). It also shows state-of-the-art performance on synthetic-to-real and ideal-to-adverse generalization.
@@
  • To the best of our knowledge, it is the first attempt to use ViT for the task of domain generalized USSS.
  • A content-enhanced mask attention mechanism is proposed enhancing the content information from image features.
  • A multi-resolution content-enhanced mask attention learning scheme is proposed for expressive content representation and style robustness.
@@
  • Extensive experiments show a large performance improvement over existing SOTA by upto 14.00% mIoU.
2 Related Work
@@
Domain Generalization has been studied on no task-specific scenarios in the field of both machine learning and computer vision. Harary et al. [13] considered domain generalization in an unsupervised manner by learning a domain bridge. Hu et al. [15] proposed a framework for image retrieval in an unsupervised setting. Zhou et al. [67] proposed a framework to generalize to new homogeneous domains. Qiao et al. [41] and Peng et al. [39] proposed to learn domain generalization from a single source domain. Many other methods have been proposed such as entropy regularization [64], casual matching [29], extrinsic-intrinsic interaction [53], balance invariance [2], batch normalization embeddings [45] and multiple latent domain modeling [30].
@@
Domain Generalized Semantic Segmentation can be regarded as the exploration of the prior unsupervised domain adaption segmentation task [38, 37, 63], but requires a larger generalization capability of a model on a variety of target domains. Existing methods focus on the generalization of in-the-wild [40], scribble [48] and multi-source images [20, 22], where substantial alterations can occur in both the content and style.
@@
Domain Generalized USSS focuses on the generalization of driving-scenes [8, 59, 33, 43, 42]. These methods use either a normalization transformation (e.g., IBN [35], IN [16], SAN [38]) or whitening transformation (e.g., IW [36], ISW [7], DIRL [57], SAW [38]) on the training domain, to enable the model to generalize better on the target domains. Other advanced methods for domain generalization in segmentation typically rely on external images to incorporate more diverse styles [23, 65, 66, 25], and leverage content consistency across multi-scale features [62]. To the best of our knowledge, all of these methods are based on CNN.
@@
Mask Transformer for Semantic Segmentation Earlier ViT based segmentation models (e.g., SegFormer [56]) follow the CNN based pipeline. Recently, the focus is on mask-level pipelines, which use the queries in the Transformer decoder to learn the masks, e.g., Segmenter [47], MaskFormer [5], Max-DeepLab [51], CMT-DeepLab [60], kMaX-DeepLab [61] and etc. More recently, Mask2Former [6] further simplified the pipeline of MaskFormer and achieves better performance.
3 Methodology
3.1 Problem Definition
@@
Domain Generalization can be formulated as a worst-case problem [24, 66, 50]. Given a source domain $\mathcal{S}$, and a set of unseen target domains $\mathcal{T}_{1},\mathcal{T}_{2},\cdots$, a model parameterized by $\theta$ with the task-specific loss $\mathcal{L}_{task}$, the generic domain generalization task can be formulated as a worst-case problem, given by
@@
Domain generalized USSS is challenging as it has similar content (e.g., semantics) but variation of
Figure 2: (a) Feature domain for urban-scenes, where the content is similar while the styles have clear gaps; (b) The goal is robust style handing and stable content representing.
style (e.g., urban landscapes, weather, season, illumination) among each domain $\mathcal{S}$, $\mathcal{T}_{1}$, $\mathcal{T}_{2}$, $\cdots$. Visual examples have been provided in Fig. 1. Here we analyze the feature domain. As shown in Fig. 2a, while the content distribution is similar, the style distribution is separated and accounts for the domain gap. Our goal is to find a representation where both style and content are similarly distributed among domains, therefore minimizing the domain gap (in Fig.2b).
3.2 Content-enhanced Mask Attention
Figure 3: (a) The proposed Content-enhanced Mask Attention (CMA). Apart from the conventional mask attention (left branch), the down-sampled feature maps, which contain more abundant content information, are used to learn content enhanced query features (right branch). (b) Framework overview of the proposed Content-enhanced Mask TransFormer (CMFormer) for domain generalized semantic segmentation. The image decoder is directly inherited from the Mask2Former [6].
Our proposed idea is to enhance the focus of the mask attention mechanism [5, 6] on content information from urban scenes. This enhancement aids the segmentation masks in concentrating on scene content while reducing sensitivity to style variations. Since content/style information is usually more pronounced in the low/high resolution features of an image, it is intuitive to amplify the influence of low-resolution features in the self-attention learning.
@@ 3.2 Content-enhanced Mask Attention
Mask Attention Mechanism learns the query features as the segmentation masks by introducing a mask attention matrix based on the self-attention mechanism. Let $\mathbf{F}_{l}\in\mathbb{R}^{(W_{l}\cdot H_{l})\times C_{F}}$ denote the image features from the image decoder with a spatial width and height of $W_{l}$ and $H_{l}$, and let $\mathbf{X}_{l}\in\mathbb{R}^{N\times C}$ denote the features of the $l^{th}$ layer in a Transformer decoder, where $N$ is the number of semantic categories. Further, $C_{F}$ and $C$ denote the channel dimensions of the image features and self-attention embeddings respectively. $\mathbf{X}_{0}$ refers to the input query features of the Transformer decoder. The key $\mathbf{K}_{l}\in\mathbb{R}^{(W_{l}\cdot H_{l})\times C}$, value $\mathbf{V}_{l}\in\mathbb{R}^{(W_{l}\cdot H_{l})\times C}$, and query $\mathbf{Q}_{l}\in\mathbb{R}^{N\times C}$ is computed by linear transformations $f_{K}$, $f_{V}$ and $f_{Q}$, respectively, given by
$\mathbf{K}_{l}=f_{K}(\mathbf{F}_{l-1}),\mathbf{V}_{l}=f_{V}(\mathbf{F}_{l-1}),\mathbf{Q}_{l}=f_{Q}(\mathbf{X}_{l-1}).$ (2)
@@
Then, the query feature $\mathbf{X}_{l}$ is computed by
@@
Content Enhancement is incorporated into the mask attention mechanism by acquiring supplementary query features from low-resolution image features. This is based on the understanding that content information tends to be more discriminative, while style information is less sensitive in low-resolution image features. To this end, a parallel branch that leverages low-resolution image features is introduced in each Transformer decoder layer (in Fig. 3a). The content-enhanced image feature $\mathbf{F}_{l}^{c}\in\mathbb{R}^{(W_{l}\cdot H_{l})\times C_{F}}$ is computed by average pooling $\rm{avgpool}$ from the original image feature $\mathbf{F}_{l}^{c}\in\mathbb{R}^{(W_{l}\cdot H_{l})\times C_{F}}$ by
$\mathbf{F}_{l}^{c}={\rm avgpool}(\mathbf{F}_{l}).$ (5)
Following Eq. 2, for the $l^{th}$ Transformer decoder, the content enhanced key $\mathbf{K}^{c}_{l}\in\mathbb{R}^{(W_{l}/2\cdot H_{l}/2)\times C}$ and value $\mathbf{V}^{c}_{l}\in\mathbb{R}^{(W_{l}/2\cdot H_{l}/2)\times C}$ are computed from the linear transformation of the down-sampled image feature $\mathbf{F}_{l}^{c}$. Then, similar to Eq. 3, the content-enhanced query feature $\mathbf{X}_{l}^{c}$ is computed as
$\mathbf{X}_{l}^{c}={\rm softmax}(\mathcal{M}_{l-1}^{c}+\mathbf{Q}_{l}\mathbf{K}_{l}^{c\mathsf{T}})\mathbf{V}_{l}^{c}+\mathbf{X}_{l-1}^{c},$ (6)
where $\mathcal{M}_{l-1}^{c}\in\{0,1\}^{N\times(H_{l}/2\cdot W_{l}/2)}$ also follows the computation of $\mathcal{M}_{l-1}$ defined in Eq. 4.
Content-aware Query Feature Fusion is proposed to merge the query $\mathbf{X}_{l}$ with content-enhanced query features $\mathbf{X}_{l}^{c}$, allowing for the highlighting of content while retaining the necessary details for semantic prediction. The fused feature $\mathbf{X}_{l}^{final}$ serves as the final output of the $l^{th}$ Transformer decoder, and it is computed as follows:
$\mathbf{X}_{l}^{final}=h_{l}([\mathbf{X}_{l},\mathbf{X}_{l}^{c}]),$ (7)
@@
Here, [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] represents the concatenation operation, and $h_{l}(\cdot)$ refers to a linear layer.
3.3 Content-enhancement for Multi-resolution Features
In the process of decoding domain-generalized segmentation predictions from coarse-to-fine image features, it is crucial to enhance the content information at each resolution. This ensures that the content remains highlighted in the query features of masks, allowing for the learning of segmentation masks that are less sensitive to style variations.
Multi-resolution Features include $\times$32, $\times$16 and $\times$8 resolution features from the image decoder, denoted as $\mathbf{F}^{\times 32}$, $\mathbf{F}^{\times 16}$ and $\mathbf{F}^{\times 8}$, respectively. Following Eq. 23, the key, value and query for $\mathbf{F}^{\times 32}$ and $\mathbf{F}^{\times 16}$ and $\mathbf{F}^{\times 8}$ are denoted by $\mathbf{K}^{\times 32}$, $\mathbf{V}^{\times 32}$, $\mathbf{Q}^{\times 32}$ and $\mathbf{K}^{\times 16}$, $\mathbf{V}^{\times 16}$, $\mathbf{Q}^{\times 16}$ and $\mathbf{K}^{\times 8}$, $\mathbf{V}^{\times 8}$, $\mathbf{Q}^{\times 8}$, respectively.
Content-enhanced Multi-resolution Features are down-sampled from $\times$32, $\times$16 and $\times$8 resolution features $\mathbf{F}^{\times 32}$ and $\mathbf{F}^{\times 16}$ and $\mathbf{F}^{\times 8}$ into $\mathbf{F}^{\times 32,c}$ and $\mathbf{F}^{\times 16,c}$ and $\mathbf{F}^{\times 8,c}$ by following Eq. 5. Also following Eq. 23, the key, value and query for $\mathbf{F}^{\times 32,c}$ and $\mathbf{F}^{\times 16,c}$ and $\mathbf{F}^{\times 8,c}$ are denoted as $\mathbf{K}^{\times 32,c}$, $\mathbf{V}^{\times 32,c}$, $\mathbf{Q}^{\times 32,c}$ and $\mathbf{K}^{\times 16,c}$, $\mathbf{V}^{\times 16,c}$, $\mathbf{Q}^{\times 16,c}$ and $\mathbf{K}^{\times 8,c}$, $\mathbf{V}^{\times 8,c}$, $\mathbf{Q}^{\times 8,c}$, respectively.
Multi-resolution Feature Fusion to the Transformer Decoder directly follows the original Mask2Former [6], which alternatively learns the image features from each different resolution. As there are 9 self-attention layers in the Transformer decoder, the first, forth, seventh layers take the ×$\times 32$ image features as input (black down-arrows in Fig. 3b), the second, fifth, eighth layers take the ×$\times 16$ image features as the input (blue downarrows in Fig. 3b), and the third, sixth, ninth layer take the ×$\times 8$ image features as the input (green downarrows in Fig. 3b).
Taking the first, forth, seventh layers with ×$\times 32$ image features as an example (here $l=1,4,7$), the content-enhanced learning process is given by
$\mathbf{X}_{l}={\rm softmax}(\mathcal{M}_{l-1}+\mathbf{Q}_{l}^{\times 32}\mathbf{K}_{l}^{\times 32\mathsf{T}})\mathbf{V}_{l}^{\times 32}+\mathbf{X}_{l-1}^{final},$ (8)
$\mathbf{X}_{l}^{c}={\rm softmax}(\mathcal{M}_{l-1}^{c}+\mathbf{Q}_{l-1}^{\times 32,c}\mathbf{K}^{\times 32,c\mathsf{T}})\mathbf{V}^{\times 32,c}+\mathbf{X}_{l-1}^{final},$ (9)
Following Eq. 7, $\mathbf{X}_{l}^{c}$ and $\mathbf{X}_{l}$ are merged to compute the final output of this decoder layer, denoted as $\mathbf{X}_{l}^{final}$. For the first layer, $\mathbf{X}_{0}^{final}$ is equivalent to $\mathbf{X}_{0}$, which is the input query features of the Transformer decoder.
The content enhance learning of the second, fifth and eighth layer follows Eq. 89 but uses the $\mathbf{K}^{\times 16,c}$, $\mathbf{V}^{\times 16,c}$, $\mathbf{Q}^{\times 16,c}$ and $\mathbf{K}^{\times 16}$, $\mathbf{V}^{\times 16}$, $\mathbf{Q}^{\times 16}$ as input. The third, sixth and ninth layer also follow Eq. 89 but takes $\mathbf{K}^{\times 8,c}$, $\mathbf{V}^{\times 8,c}$, $\mathbf{Q}^{\times 8,c}$ and $\mathbf{K}^{\times 8}$, $\mathbf{V}^{\times 8}$, $\mathbf{Q}^{\times 8}$ as input.
3.4 Network Architecture and Implementation Details
@@
The overall framework is shown in Fig. 3b. The Swin-Base Transformer [27] is used as the backbone, and the image decoder is inherited from Mask2Former [6]. The pre-trained model from ImageNet [9] is utilized for initialization.

### v4

@@ Abstract
In this paper, we propose a Content-enhanced Mask TransFormer (CMFormer) for domain-generalized USSS. The main idea is to enhance the focus of the fundamental component, the mask attention mechanism, in Transformer segmentation models on content information. We have observed through empirical analysis that a mask representation effectively captures pixel segments, albeit with reduced robustness to style variations. Conversely, its lower-resolution counterpart exhibits greater ability to accommodate style variations, while being less proficient in representing pixel segments. To harness the synergistic attributes of these two approaches, we introduce a novel content-enhanced mask attention mechanism. It learns mask queries from both the image feature and its down-sampled counterpart, aiming to simultaneously encapsulate the content and address stylistic variations. These features are fused into a Transformer decoder and integrated into a multi-resolution content-enhanced mask attention learning scheme.
@@
Extensive experiments conducted on various domain-generalized urban-scene segmentation datasets demonstrate that the proposed CMFormer significantly outperforms existing CNN-based methods for domain-generalized semantic segmentation, achieving improvements of up to 14.00% in terms of mIoU (mean intersection over union). The source code is publicly available at https://github.com/BiQiWHU/CMFormer.
Introduction
@@
Urban-scene semantic segmentation (USSS) is a challenging problem because of the large scene variations due to changing landscape, weather, and lighting conditions (Sakaridis, Dai, and Van Gool 2021; Mirza et al. 2022; Diaz-Ruiz et al. 2022; Chen et al. 2022). Unreliable USSS can pose a significant risk to road users. Nevertheless, a segmentation model trained on a specific dataset cannot encompass all urban scenes across the globe. As a result, the segmentation model is prone to encountering unfamiliar urban scenes during the inference stage. Hence, domain generalization is essential for robust USSS (Pan et al. 2018; Huang et al. 2019a; Choi et al. 2021), where a segmentation model can effectively extrapolate its performance to urban scenes that it hasn’t encountered before (Fig. 1). In contrast to common domain generalization, domain generalized USSS requires special attention because the domain gap is mainly caused by large style variations whereas changes in semantics largely remain consistent (e.g. as shown in the example in Fig. 2).
Figure 1: (a) In domain generalized USSS, the domain gap is mainly from the extremely-varied styles. (b) A segmentation model is supposed to show good generalization on unseen target domains.
@@
Existing approaches can be divided into two groups. One group focuses on the style de-coupling. This is usually achieved by a normalization (Pan et al. 2018; Huang et al. 2019a; Peng et al. 2022) or whitening (Pan et al. 2019; Choi et al. 2021; Xu et al. 2022; Peng et al. 2022) transformation. However, the de-coupling methodology falls short as the content is not learnt in a robust way. The other group is based on adverse domain training (Zhao et al. 2022; Lee et al. 2022; Zhong et al. 2022). However, these methods usually do not particularly focus on urban styles and therefore their performance is limited.
Recent work has shown that mask-level segmentation Transformer (e.g., Mask2Former) (Ding et al. 2023) is a scalable learner for domain generalized semantic segmentation. However, based on our empirical observations, a high-resolution mask-level representation excels at capturing content down to pixel semantics but is more susceptible to style variations. Conversely, its down-sampled counterpart is less proficient in representing content down to pixel semantics but exhibits greater resilience to style variations.
Figure 2: Domain-generalized USSS demonstrates a distinctive feature of consistent content with diverse styles. An example is given for BDD100K and GTA5.
@@
A novel Content-enhanced mask attention (CMA) mechanism is proposed. It jointly leverages both mask representation and its down-sampled counterpart, which show complementary properties on content representing and handling style variation. Jointly using both features helps the style to be uniformly distributed while the content to be stabilized in a certain cluster. The proposed CMA takes the original image feature together with its down-sampled counterpart as input. Both features are fused to learn a more robust content from their complementary properties.
The proposed Content-enhanced mask attention (CSMA) mechanism can be integrated into existing mask-level segmentation Transformer in a learnable fashion. It consists of three key steps, namely, exploiting high-resolution properties, exploiting low-resolution properties, and content-enhanced fusion. Besides, it can also be seamlessly adapted to multi-resolution features. A novel Content-enhanced Mask TransFormer (CMFormer) is proposed for domain-generalized USSS.
@@
Large-scale experiments are conducted with various domain generalized USSS settings, i.e., trained on one dataset from (Richter et al. 2016; Ros et al. 2016; Cordts et al. 2016; Neuhold et al. 2017; Yu et al. 2018) as the source domain, and validated on the rest of the four datasets as the unseen target domains. All the datasets contain the same 19 semantic categories as the content, but vary in terms of scene styles. The experiments show that the proposed CMFormer achieves upto 14.00% mIoU improvement compared to the state-of-the-art CNN based methods (e.g., SAW (Peng et al. 2022), WildNet (Lee et al. 2022)). Furthermore, it demonstrates a mIoU improvement of up to 1.7% compared to the modern HGFormer model(Ding et al. 2023). It also shows state-of-the-art performance on synthetic-to-real and clear-to-adverse generalization.
@@
  • A Content-enhanced mask attention (CMA) mechanism is proposed to leverage the complementary Content and style properties from mask-level representation and its down-sampled counterpart.
  • On top of CMA, a Content-enhanced Mask Transformer (CMFormer) is proposed for domain generalized urban-scene semantic segmentation.
@@
  • Extensive experiments show a large performance improvement over existing SOTA by upto 14.00% mIoU, and HGFormer by upto 1.7% mIoU.
Related Work
Domain Generalization
@@
has been studied on no task-specific scenarios in the field of both machine learning and computer vision. Hu et al. (Hu and Lee 2022) proposed a framework for image retrieval in an unsupervised setting. Zhou et al. (Zhou et al. 2020) proposed a framework to generalize to new homogeneous domains. Qiao et al. (Qiao, Zhao, and Peng 2020) and Peng et al. (Peng, Qiao, and Zhao 2022) proposed to learn domain generalization from a single source domain. Many other methods have also been proposed (Zhao et al. 2020; Mahajan, Tople, and Sharma 2021; Wang et al. 2020; Chattopadhyay, Balaji, and Hoffman 2020; Segu, Tonioni, and Tombari 2023).
Domain Generalized Semantic Segmentation
@@
can be regarded as the exploration of unsupervised domain adaption segmentation (Peng et al. 2022, 2021; Yue et al. 2019a), but requires a larger generalization capability of a model on a variety of target domains. Existing methods focus on the generalization of in-the-wild (Piva, de Geus, and Dubbelman 2023), scribble (Tjio et al. 2022) and multi-source images (Kim et al. 2022; Lambert et al. 2020), where substantial alterations can occur in both the content and style.
Domain Generalized USSS
@@
focuses on the generalization of driving-scenes (Cordts et al. 2016; Yu et al. 2018; Neuhold et al. 2017; Ros et al. 2016; Richter et al. 2016). These methods use either a normalization transformation (e.g., IBN (Pan et al. 2018), IN (Huang et al. 2019a), SAN (Peng et al. 2022)) or whitening transformation (e.g., IW (Pan et al. 2019), ISW (Choi et al. 2021), DIRL (Xu et al. 2022), SAW (Peng et al. 2022)) on the training domain, to enable the model to generalize better on the target domains. Other advanced methods for domain generalization in segmentation typically rely on external images to incorporate more diverse styles (Lee et al. 2022; Zhao et al. 2022; Zhong et al. 2022; Li et al. 2023), and leverage content consistency across multi-scale features (Yue et al. 2019b). To the best of our knowledge, all of these methods are based on CNN.
@@
Mask Transformer for Semantic Segmentation uses the queries in the Transformer decoder to learn the masks, e.g., Segmenter (Strudel et al. 2021), MaskFormer (Cheng, Schwing, and Kirillov 2021). More recently, Mask2Former (Cheng et al. 2022) further simplified the pipeline of MaskFormer and achieves better performance.
Preliminary
Problem Definition
@@
Domain Generalization can be formulated as a worst-case problem (Li, Namkoong, and Xia 2021; Zhong et al. 2022; Volpi et al. 2018). Given a source domain $\mathcal{S}$, and a set of unseen target domains $\mathcal{T}_{1},\mathcal{T}_{2},\cdots$, a model parameterized by $\theta$ with the task-specific loss $\mathcal{L}_{task}$, the generic domain generalization task can be formulated as a worst-case problem, given by
@@
As summarized in ISW (Choi et al. 2021), domain generalized USSS is challenging as it has similar content (e.g., semantics) but may vary in style (e.g., urban landscapes, weather, season, illumination) among each domain.
Similar content. In Fig. 2a, samples from BDD100K and from GTA5 have same scene categories, namely, tree, cars, road and building. Without loss of generality, samples from all the five USSS datasets contain the same 19 scene-object categories. These scene-object categories all commonly exist in urban-scenes, either seen or unseen during training.
Varied styles. In Fig. 2a, we present an example sourced from BDD100K, showcasing a rainy, daytime scene from a European street, among other factors. In contrast, the GTA5 sample embodies a clear, nighttime scene from an American street, along with other attributes. Without compromising the breadth of representation, samples across all five USSS datasets exhibit an even broader spectrum of stylistic variations. These variations span weather conditions such as clear, rainy, and snowy scenarios, while also encompassing different times of day—CityScapes predominantly features daytime scenes, whereas GTA5 portrays scenes from both day and night. Regrettably, despite the pronounced diversity in style, this extensive variation may still not encompass the entirety of potential conditions in urban environments.
Content-style Feature Space
Here we analyze the feature space. Figure 3a illustrates that in the context of domain-generalized USSS, samples from distinct domains might exhibit analogous patterns and cluster tightly along the content dimension. Conversely, samples from diverse domains may segregate into separate clusters along the style dimension.
Figure 3: In the domain-generalized USSS setting, within the content-style space, samples from various domains tend to cluster closely along the content dimension while displaying dispersion along the style dimension. (b) An optimal generalized semantic segmentation scenario would involve uniform distribution of styles while maintaining content stability (as indicated by the brown bounding box).
An optimal and adaptable segmentation representation should achieve content stability while simultaneously exhibiting resilience in the face of significant style variations. Illustrated in Figure 3b, our objective is to cultivate a content-style space wherein: 1) samples from diverse domains can occupy analogous positions along the content dimension, and 2) samples can be uniformly dispersed across the style dimension. Both learning objectives allow us to therefore minimize the domain gap.
Overall Idea
Recent work has shown that mask-level segmentation Transformer (e.g., Mask2Former) (Ding et al. 2023) is a scalable learner for domain generalized semantic segmentation. Unfortunately, we empirically observed that, a mask-level representation is better at representing content, but more sensitive to style variations (similar to Fig. 3a); its low-resolution counterpart, on the contrary, is less capable to represent content, but more robust to the style variations (similar to the style dimensions in Fig. 3b).
Overall, the mask representation and its down-sampled counterpart shows complementary properties when handling samples from different domains. Thus, it is natural to jointly leverage both mask representation and its down-sampled counterparts, so as to at the same time stabilizing the content and be insensitive to the style variation.
Difference between Existing Pipelines
Existing methods usually focus on decoupling the styles from urban scenes, so that along the style dimension the samples from different domains are more uniformly distributed.
In contrast, the proposed method intends to leverage the content representation ability of mask-level features (content dimension of Fig. 3a) and the style handling ability of its down-sampled counterpart (style dimension of Fig. 3b), so as to realize the same learning objective.
Methodology
Recap on Mask Attention
Recent studies show that the mask-level pipelines (Strudel et al. 2021; Cheng, Schwing, and Kirillov 2021; Cheng et al. 2022) have stronger representation ability than conventional pixel-wise pipelines for semantic segmentation, which can be attributed to the mask attention mechanism.
@@
It learns the query features as the segmentation masks by introducing a mask attention matrix based on the self-attention mechanism. Let $\mathbf{F}_{l}$, $\mathbf{X}_{l}$ denote the image features from the image decoder, and the features of the $l^{th}$ layer in a Transformer decoder, respectively. When $l=0$, $\mathbf{X}_{0}$ refers to the input query features of the Transformer decoder.
@@
The key $\mathbf{K}_{l}$ and value $\mathbf{V}_{l}$ on $\mathbf{F}_{l-1}$ are computed by linear transformations $f_{K}$ and $f_{V}$, respectively. And the query $\mathbf{Q}_{l}$ on $\mathbf{X}_{l-1}$ is computed by linear transformation $f_{Q}$. Then, the query feature $\mathbf{X}_{l}$ is computed by
@@
Exploiting High-Resolution Properties
Figure 4: (a) The proposed Content-enhanced Mask Attention (CMA) consists of three key steps, namely, exploiting high-resolution properties (in green), exploiting low-resolution properties (in brown), and content enhanced fusion (in gray). (b) Framework overview (in yellow) of the proposed Content-enhanced Mask TransFormer (CMFormer) for domain generalized semantic segmentation. The image decoder is directly inherited from the Mask2Former (Cheng et al. 2022).
Highlighted within the green block in Figure 4, our empirical observations reveal that the high-resolution mask representation exhibits the following characteristics: 1) greater proficiency in content representation, and 2) reduced susceptibility to domain variation. However, achieving uniform mixing of samples from four domains presents a challenge.
To leverage the properties from high-resolution mask representations, we use the self-attention mechanism to exploit the amplified content representation from $\mathbf{X}_{l}$. Let $\mathbf{Q}_{\mathbf{X}_{l}}$, $\mathbf{V}_{\mathbf{X}_{l}}$ and $\mathbf{K}_{\mathbf{X}_{l}}$ denote its query, value and key. Then, the self-attention is computed as
${\rm Attention}(\mathbf{Q}_{\mathbf{X}_{l}},\mathbf{K}_{\mathbf{X}_{l}},\mathbf{V}_{\mathbf{X}_{l}})={\rm Softmax}(\frac{\mathbf{Q}_{\mathbf{X}_{l}}\mathbf{K}_{\mathbf{X}_{l}}}{\sqrt{d_{k}}})\mathbf{V}_{\mathbf{X}_{l}},$ (4)
where $\rm{Softmax}$ denotes the softmax normalization function, and the final output is denoted as $\tilde{\mathbf{X}}_{l}$.
Exploiting Low-Resolution Properties
As shown in the brown block of Fig. 4, the low-resolution mask-level representation has the following properties: 1) less qualified to represent the content; 2) more capable to handle the style variation. In the feature space, samples from different domains are more uniformly distributed. We propose to build a low-resolution mask representation derived from its high-resolution counterpart. This approach capitalizes on the attributes of the low-resolution representation to effectively address domain variations.
The low-resolution counterpart $\mathbf{F}_{l}^{d}$ is computed by average pooling $\rm{avgpool}$ from the original image feature $\mathbf{F}_{l}$ by
$\mathbf{F}_{l}^{d}={\rm avgpool}(\mathbf{F}_{l}),$ (5)
where the width and height of $\mathbf{F}_{l}$ is both twice the width and height of $\mathbf{F}_{l}^{d}$.
Similarly, the key and value from $\mathbf{F}_{l}^{d}$ are computed by linear transformations, and can be denoted as $\mathbf{K}^{d}_{l}$ and $\mathbf{V}^{d}_{l}$, respectively. The query from $\mathbf{X}_{l-1}^{d}$ is also computed by linear transformation, and can be denoted as $\mathbf{K}^{d}_{l}$. The mask attention on the low-resolution feature $\mathbf{X}_{l}^{d}$ is computed as
$\mathbf{X}_{l}^{d}={\rm softmax}(\mathcal{M}_{l-1}^{d}+\mathbf{Q}_{l}\mathbf{K}_{l}^{d\mathsf{T}})\mathbf{V}_{l}^{d}+\mathbf{X}_{l-1}^{d}.$ (6)
To exploit the properties from the low-resolution mask representation $\mathbf{X}_{l}^{d}$, we use the self-attention mechanism. Let $\mathbf{Q}_{\mathbf{X}_{l}^{d}}$, $\mathbf{V}_{\mathbf{X}_{l}^{d}}$ and $\mathbf{K}_{\mathbf{X}_{l}^{d}}$ denote its query, value and keys. Then, the self-attention is computed by
${\rm Attention}(\mathbf{Q}_{\mathbf{X}_{l}^{d}},\mathbf{K}_{\mathbf{X}_{l}^{d}},\mathbf{V}_{\mathbf{X}_{l}^{d}})={\rm Softmax}(\frac{\mathbf{Q}_{\mathbf{X}_{l}^{d}}\mathbf{K}_{\mathbf{X}_{l}^{d}}}{\sqrt{d_{k}}})\mathbf{V}_{\mathbf{X}_{l}^{d}}.$ (7)
The final output is denoted as $\tilde{\mathbf{X}}_{l}^{d}$. It inherits the characteristics of the low-resolution mask representation, which is adept at accommodating style variations while being less resilient in capturing pixel-level intricacies.
Content-enhanced Fusion
Our idea is to leverage the complementary properties of mask-level representation and its down-sample counterpart, so as to enhance both the pixel-wise representing and style variation handing (shown in the gray box of Fig. 4). The joint use of both representations aids the segmentation masks in concentrating on scene content while reducing sensitivity to style variations.
To this end, we fuse both representations $\tilde{\mathbf{X}}_{l}$ and $\tilde{\mathbf{X}}_{l}^{d}$ in a simple and straight-forward way. The fused feature $\mathbf{X}_{l}^{final}$ serves as the final output of the $l^{th}$ Transformer decoder, and it is computed as follows:
$\mathbf{X}_{l}^{final}=h_{l}([\tilde{\mathbf{X}}_{l},\tilde{\mathbf{X}}_{l}^{d}]),$ (8)
@@
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] represents the concatenation operation, and $h_{l}(\cdot)$ refers to a linear layer.
Network Architecture and Implementation Details
@@
The overall framework is shown in the yellow box of Fig. 4. The Swin Transformer (Liu et al. 2021) is used as the backbone. The pre-trained backbone from ImageNet (Deng et al. 2009) is utilized for initialization.